OpenStudy (aravindg):

The sum of 2 non-zero nos is 8.Then min value of the sum of their reciprcocals is:

OpenStudy (aravindg):

@amistre64 , @asnaseer , @TuringTest , @experimentX

OpenStudy (aravindg):

@Ishaan94 , @satellite73

OpenStudy (asnaseer):

\[x+y=8\]and we want to minimize:\[\frac{1}{x}+\frac{1}{y}\]

OpenStudy (aravindg):

ys

OpenStudy (asnaseer):

so just find x interms of y from first equation. substitute into second and use calculas to find minimum.

OpenStudy (aravindg):

by equating dy/dx=0 ryt?

OpenStudy (asnaseer):

yes - remember to also consider value of \(\frac{d^2y}{dx^2}\) to determine if its a min, max or point of inflection

OpenStudy (aravindg):

k